OpenStudy (anonymous):

Factor: x2 + 15x + 54 ?

OpenStudy (anonymous):

do u know how to do c b

OpenStudy (anonymous):

and yo ufind the x values

OpenStudy (anonymous):

x intercepts

OpenStudy (is3535):

Step by step solution : Step 1 : Raise x to the 2nd power Exponentiaion : Equation at the end of step 1 : (x2 - 15x) - 54 Step 2 : Simplify x2-15x - 54 Trying to factor by splitting the middle term 2.1 Factoring x2-15x-54 The first term is, x2 its coefficient is 1 . The middle term is, -15x its coefficient is -15 . The last term, "the constant", is -54 Step-1 : Multiply the coefficient of the first term by the constant 1 • -54 = -54 Step-2 : Find two factors of -54 whose sum equals the coefficient of the middle term, which is -15 . -54 + 1 = -53 -27 + 2 = -25 -18 + 3 = -15 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -18 and 3 x2 - 18x + 3x - 54 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-18) Add up the last 2 terms, pulling out common factors : 3 • (x-18) Step-5 : Add up the four terms of step 4 : (x+3) • (x-18) Which is the desired factorization Final result : (x + 3) • (x - 18)

OpenStudy (anonymous):

A. (x + 9)(x + 6) B. (x - 9)(x - 6) C. (x + 10)(x + 5) D. (x + 8)(x + 7)

OpenStudy (anonymous):

those are my choices

OpenStudy (anonymous):